OpenStudy (anonymous):

x+3x+6x-20=180 x=10 how do I get 10 for x?

OpenStudy (anonymous):

collect like terms on the left to get \[ 10x-20=180\] then add 20 to both sides to get \[10x=200\] so \[x=20\] not 10

OpenStudy (anonymous):

in my math lab for class it gives me the answer x=10 but I can figure out how they came up with 10?? it takes me to the next step being 10x-20=180

OpenStudy (anonymous):

oh i see, you are not asking for x, just what is \[x+3x+6x\] which is \[10x\] for the same reason that a bottle of beer plus 3 bottles of beer plus 6 bottles of beer is ten bottles of beer

OpenStudy (anonymous):

OHH I got it now because x is alone in the equation it equals one so x+3+6=10 right?

OpenStudy (anonymous):

thanks

Directrix (directrix):

Think of the "alone" x as 1x. Then, x+3x+6x-20=180 1x + 3x + 6x - 20 = 180 10x = 200